OpenStudy (anonymous):

Michael and Kathryn bowl together and their combined total score for one game was 425 points. Michael’s score was 70 less than twice Kathryn’s. What were their scores? Which is a system of equations to model the problem if x represents Kathryn’s score and y represents Michael’s score? x + y = 70 y = 2x – 425 x – y = 425 y = 2x – 70 x + y = 425 y = 2x – 70 x + y = 425 y = 2x + 70

OpenStudy (kendricklamar2014):

I am guessing those are the answer choices. What is your answer @WildQueen

OpenStudy (anonymous):

ig it is B

OpenStudy (kendricklamar2014):

B is actually wrong. This is why: Let x = Kathryn and y = Micheal \[425 = x + y\] \[y = 2x - 70\]

OpenStudy (kendricklamar2014):

Just switch it around to get: \[x + y = 425\] \[y = 2x – 70\]

OpenStudy (kendricklamar2014):

Thats the answer ^_^